40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Aug 5 11:49:49 2009, Stephanie, Update, General, Multiply Resonant EOM Update OSASetup.pngSMAPic.pngWithSMA.png
    Reply  Wed Aug 5 15:18:12 2009, Stephanie, Update, General, Multiply Resonant EOM Update SMA.png
Message ID: 1834     Entry time: Wed Aug 5 11:49:49 2009     Reply to this: 1835
Author: Stephanie 
Type: Update 
Category: General 
Subject: Multiply Resonant EOM Update 

I have spent the past couple of days gathering optics and mounts so that I can observe the modulation of the EOM attached to the circuit I built using the optical spectrum analyzer (OSA). A rough diagram of the planned layout is attached.

I also built a short SMA cable so that the EOM did not have to be connected directly to the circuit box. The cable is shown attached to the EOM and circuit box in the attached photo. After checking to make sure that all of the connections in the cable were sound, I remeasured the input impedance of the circuit; the impedance measurement (black) is shown in the attached plot with the impedance before the SMA cable was added with and without the box (green and blue, respectively--these two are almost identical). The new impedance has a strange shape compared to the original measurements; I'd like to understand this a little better, since adding extra inductance in LTSpice doesn't seem to have that effect. Since I had already taken apart the setup used for the previous impedance measurements, I had to rebuild and recalibrate the setup; I guess the difference could be something about the new calibration, but I don't really think that that's the case.

Attachment 1: OSASetup.png  27 kB  | Hide | Hide all
OSASetup.png
Attachment 2: SMAPic.png  459 kB  | Hide | Hide all
SMAPic.png
Attachment 3: WithSMA.png  23 kB  | Hide | Hide all
WithSMA.png
ELOG V3.1.3-